SM.61 A small but growing online retailer, Nile Corporation, has shown impressive growth in sales over the past several years, with sales this past year at $836,000.

If the company has a net profit margin of 2.75 percent, what would its net profit be (in dollars)? (Display your answer as a whole number.)

If in the next year the company achieves its revenue growth target of 13 percent, what would its total revenue be? (Display your answer as a whole number.)

If in the next year the company achieves its revenue growth target of 13 percent, and assuming its profit margin remained unchanged at 2.75 percent, what would its total profit be for next year? (Display your answer as a whole number.)

If the company achieves its revenue growth target of 13 percent, by how many dollars will revenue increase? (Display your answer as a whole number.)

If the company achieves its revenue growth target of 13 percent, by how many dollars will net profit increase? (Display your answer as a whole number.)

Using the original revenue number of $836,000, if the company spends 64 percent of its revenue on purchases, what would be its purchasing expense? (Display your answer as a whole number.)

Assuming that revenues stayed flat (meaning the company did not try to increase sales by the 13 percent target), by what percentage would they have to decrease purchasing expenses to equal the increased profit that would have come from a 13 percent increase to revenues? (Write your answer as a percentage, and display your answer to two decimal places.)
Answer #1

Note :- As per Chegg rules answering the first 4 questions

Given, Sales in the Past Year = $836,000

Net Profit Margin = 2.75%

We know, Net Profit Margin % = (Net Profit / sales)*100

So Net Profit = $(836,000*2.75%) = $22,990

If in the next year 13% increase in the revenue growth is achieved by the company.

Total revenue would be = Previous revenue * 113%

                                    = $(836,000 * 1.13)     = $944,680

Now if the company will achieve a revenue growth of 2.75% and its profit margin has remain unchanged at 2.75% (as of this year)

Hence the total net profit for next year = $(944,680*2.75%) = $25,978.7

If the 13% margin is achieved then revenue in terms of Dollar will increase for the company by the value = (Next Year Revenue – Previous Year Revenue)

          = $ (944,680 - 836,000)   = $108,680